OpenStudy (anonymous):

If h(x)=x^2-5, find h(-3) A.-14 B.-11 C.4

OpenStudy (anonymous):

plug -3 into h(x) to get ur answer

OpenStudy (luigi0210):

f(-3)=(-3)^2-5

OpenStudy (anonymous):

I dont understand how to do the problem

OpenStudy (rulnick):

"h(-3)" means "h of -3" (not "h times -3"), meaning the function h evaluated at x=-3. So you take h(x) and replace x with -3: h(-3) = (-3)^2-5 = 9 - 5 = ... Do you see?

OpenStudy (anonymous):

Oh ok thank you. c:

OpenStudy (rulnick):

welcome